### v2

# Analysis of the $q\bar{q}\to Z^{*}\to hA\to 4\tau$ process within the lepton-specific 2HDM at the LHC

Abstract
We analyse light Higgs scalar and pseudoscalar associated hadro-production in the 2-Higgs Doublet Model (2HDM) Type-X (or lepton-specific) within the parameter space allowed by theoretical self-consistency requirements as well as the latest experimental constraints from the Large Hadron Collider (LHC), precision data and $B$ physics. Over the viable regions of such a scenario, the Standard Model-like Higgs boson discovered at the LHC in 2012 is the heavier CP-even state $H$. Furthermore, in the Type-X scenario, due to large $\tan\beta$, the lighter Higgs scalar $h$ and the pseudoscalar $A$ mainly decay into two $\tau$ leptons. Therefore, we concentrate on analysing the signal process $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}\to\ell\nu_{\ell}\ell\nu_{\ell}\tau_{h}\tau_{h}$ (where $\ell=e,\mu$ whereas $\tau_{h}$ represents the hadronic decay of the $\tau$) and explore the feasibility of conducting such a search at the LHC with a centre-of-mass energy of $\sqrt{s}~{}=~{}\text{14 TeV}$ and a luminosity of $L~{}=~{}\text{300}~{}\text{fb}^{-1}$. To suppress the huge SM background, we confine ourselves to consider the fraction of signal events with two same-sign $\tau$ leptons further decaying into same-sign leptons while the other two $\tau$ leptons decay hadronically. We find that a combination of kinematical selection and machine learning (ML) analysis will yields significant sensitivity to this process at the end of the LHC Run 3.
KEYWORDS: 2-Higgs Doublet Model, Type-X Scenario, Higgs Physics
I Introduction
The LHC has played an indispensable role in improving our understanding of the microcosmic subatomic structure, particularly with the discovery of a 125 GeV scalar particle by the ATLAS and CMS experimental teams in 2012 Aad:2012tfa ; Chatrchyan:2012ufa . This was the last particle predicted by the Standard Model (SM) and its properties measured in the above experiments are currently consistent with the SM prediction at the $5~{}\sigma$ level ATLAS:2022vkf ; CMS:2022dwd . Investigating the Higgs sector at the LHC, where the presence of additional (pseudo)scalars is possible, remains an intriguing task, particularly in new physics models with additional doublet fields. Therefore, to discover extra Higgs bosons at the LHC is one of the prime targets of Beyond the SM (BSM) signal searches.
One of the simplest extensions of the SM is the 2HDM, which contains two complex Higgs doublets Lee:1973iz ; Deshpande:1977rw ; Branco:2011iw . After Electro-Weak Symmetry Breaking (EWSB), there are three Goldstone bosons, which are ‘eaten’ by the $W^{\pm}$ and ${Z}$ bosons while the remaining five degrees of freedom incarnate into physical Higgs bosons: two neutral CP-even scalars (${h}$ and ${H}$, with $M_{{h}}\textless M_{{H}}$), one CP-odd pseudoscalar (${A}$) and two charged Higgs states ${H}^{\pm}$ (with mixed CP status).
In order to be consistent with the stringent experimental constraints from Flavour Changing Neutral Currents (FCNCs) at the tree level, typically, a $Z_{2}$ symmetry is introduced into the Yukawa sector so that each type of fermion only couples to one of the doublets of the 2HDM Glashow:1976nt . Based on the $Z_{2}$ charge assignments of the Higgs doublets, we can define four basic scenarios, known as (Yukawa) types. In the lepton-specific scenario (or Type-X), the additional Higgs doublet $\phi_{2}$ only couples to the lepton sector and not to the quark sector, where the light Higgs boson mainly decays to $\tau^{+}\tau^{-}$ within a wide range of parameter spaces Aoki:2009ha .
Motivated by recent searches for $\tau^{+}\tau^{-}\tau^{+}\tau^{-}$ by the CMS experiment CMS:2015twz ; CMS:2017dmg ; CMS:2019spf , we will study the process $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}$ in the 2HDM Type-X. In this scenario, the $h/A\to\tau^{+}\tau^{-}$ decay could be enhanced, resulting in a sizeable $4\tau$ final state. Therefore, it is worthwhile to examine whether this statement is robust enough after taking into account the effects of parton shower, hadronisation, heavy flavour decays and detector resolution.
The recent results by the (muon) $g-2$ experiment conducted by the Fermilab National Accelerator Laboratory (FNAL), when combined with the old ones from the Brookhaven National Laboratory (BNL), yield a 4.2 $\sigma$ discrepancy between the experimental data and the theoretical prediction from the SM:
$\Delta a_{\mu}=a_{\mu}^{\exp}-a_{\mu}^{\rm SM}=(25.1\pm 5.9)\times 10^{10},$ (1)
where $a_{\mu}=(g-2)_{\mu}/2$  Albahri2021 . One can explain this using 2HDM Type-X model, though, since the couplings of the new Higgs bosons to the SM fermions are proportional to $\text{tan}\beta$, thereby enhancing the corresponding contributions to $\Delta a_{\mu}$ Jueid2023 ; Iguro:2023tbk . Although a recent measurement has already shown possibilities of restricting the parameter spaces of the Type-X model in the large $\text{tan}\beta$ region CMS:CMS23007 , further research is still required to be performed.
In this paper, we assume that the heavier CP-even Higgs boson $H$ is the observed SM-like Higgs boson whose properties agree with the LHC measurements. We then perform a detailed Monte Carlo (MC) study on the signal process $pp\to Z^{*}\to hA\to 4\tau\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$ (with $l=e,\mu$ and $\tau_{h}$ representing hadronic $\tau$-decays) and relevant background processes, to examine its feasibility at the LHC. We do so after taking into account relevant theoretical and experimental conditions. In order to improve the experimental significance, we will require two Same-Sign (SS) leptons. Our analysis demonstrates that, by using this signature, the relevant parameter space can be either discovered or ruled out by using the current integrated luminosity at the LHC.
The paper is organised as follows. In the next section, we briefly describe the 2HDM and its Yukawa scenarios, then introduce a few Benchmark Points (BPs) for our MC analysis which pass all present theoretical and experimental constraints. In the following section, we perform a detailed collider analysis of these BPs and examine the potential to discover the aforementioned signature within the 2HDM Type-X scenario. Finally, we present some conclusions.
II The 2HDM
The Higgs sector of the 2HDM consists of two weak isospin doublets with hypercharge $Y=1$. The most general Higgs potential for the 2HDM that complies with the ${SU(2)_{L}\times U(1)_{Y}}$ gauge structure of the EW sector of the SM has the following form Branco:2011iw :
[…]
$+\frac{1}{2}\lambda_{1}(\phi_{1}^{\dagger}\phi_{1})^{2}+\frac{1}{2}\lambda_{2}(\phi_{2}^{\dagger}\phi_{2})^{2}+\lambda_{3}(\phi_{1}^{\dagger}\phi_{1})(\phi_{2}^{\dagger}\phi_{2})+\lambda_{4}(\phi_{1}^{\dagger}\phi_{2})(\phi_{2}^{\dagger}\phi_{1})~{}$
$+\frac{1}{2}\left[\lambda_{5}(\phi_{1}^{\dagger}\phi_{2})^{2}+\rm{h.c.}\right]+~{}\left\{\left[\lambda_{6}(\phi_{1}^{\dagger}\phi_{1})+\lambda_{7}(\phi_{2}^{\dagger}\phi_{2})\right](\phi_{1}^{\dagger}\phi_{2})+\rm{h.c.}\right\},$
where $\phi_{1}$ and $\phi_{2}$ are the two Higgs doublet fields. By hermiticity of such a potential, $\lambda_{1,2,3,4}$ as well as $m_{11,22}^{2}$ are real parameters, while $\lambda_{5,6,7}$ and $m_{12}^{2}$ can be complex, in turn enabling possible Charge and Parity (CP) violation effects in the Higgs sector. Upon enforcing two minimisation conditions of the potential, $m^{2}_{11}$ and $m^{2}_{22}$ can be replaced by $v_{1,2}$, which are the Vacuum Expectation Values (VEV) of the Higgs doublets $\phi_{1,2}$, respectively. Moreover, the coupling $\lambda_{1,2,3,4,5}$ can be substituted by the four physical Higgs masses (i.e., $M_{{h}},M_{{H}},M_{{A}}$ and $M_{{H}^{\pm}}$) and the parameter $\text{sin}(\beta-\alpha)$, where $\alpha$ and $\beta$ are, respectively, the mixing angles between the CP-even states and the angle related to the VEV values, i.e., $\tan\beta=v_{1}/v_{2}$. Thus, independent input parameters can be taken as $M_{{h}},M_{{H}},M_{{A}}$ and $M_{{H}^{\pm}}$, $\lambda_{6}$, $\lambda_{7}$, $\text{sin}(\beta-\alpha)$, $\text{tan}\beta$ and $m^{2}_{12}$.
If both Higgs doublet fields of the general 2HDM couple to all fermions, the ensuing scenario can induce FCNCs in the Yukawa sector at tree level. As intimated, to remedy this, a $Z_{2}$ symmetry is imposed on the Lagrangian such that each fermion type interacts with only one of the Higgs doublets Glashow:1976nt . As a consequence, there are four possible types of 2HDM, namely Type-I, Type-II, Type-X (or lepton-specific) and Type-Y (or flipped). However, such a symmetry is explicitly broken by the quartic couplings $\lambda_{6,7}$ and softly broken by the (squared) mass term $m^{2}_{12}$. In what follows, we shall consider a CP-conserving (i.e., $m^{2}_{12}$ and $\lambda_{5}$ are real) 2HDM Type-X and assume $\lambda_{6}=\lambda_{7}=0$ to forbid explicit breaking of $Z_{2}$ while also taking $m^{2}_{12}$ to be generally small, thereby preventing large FCNCs at tree level, which are incompatible with experiment.
In general, the couplings of the neutral and charged Higgs bosons to fermions can be described by the Yukawa Lagrangian given by Branco:2011iw
$-{\mathcal{L}}_{\rm Yukawa}=\sum_{f=u,d,l}\left(\frac{m_{f}}{v}\kappa_{f}^{h}\bar{f}fh+\frac{m_{f}}{v}\kappa_{f}^{H}\bar{f}fH-i\frac{m_{f}}{v}\kappa_{f}^{A}\bar{f}\gamma_{5}fA\right)+$
[…]
where $\kappa_{f}^{S}$ ($S=h,H$ and $A$) are the Yukawa couplings of the frmion $f$ in the 2HDM, which are illustrated in Tab. 1 for the Type-X under consideration. Here, $V_{ud}$ refers to a Cabibbo-Kobayashi-Maskawa (CKM) matrix element and $P_{L,R}$ denote the left- and right-handed projection operators, respectively. The coupling of the two CP-even states $h$ and $H$ to the gauge bosons $VV$ ($V=W^{\pm},Z$) are proportional to $\sin(\beta-\alpha)$ and $\cos(\beta-\alpha)$, respectively. If we assume that either $h$ or $H$ can be the observed SM-like Higgs boson, the coupling to gauge bosons is obtained for $h$ when $\cos(\beta-\alpha)\rightarrow 0$ and for $H$ when $\sin(\beta-\alpha)\rightarrow 0$. Therefore, each scenario can explain the 125 GeV Higgs signal at the LHC. Following our works Arhrib:2021xmc ; Wang:2021pxc ; Arhrib:2021yqf ; Li:2023btx , we shall focus in the present paper on the scenario where $H$ mimics the observed signal with mass  $\sim\,125$ GeV (as previously intimated).
$\kappa_{u}^{S}$ $\kappa_{d}^{S}$ $\kappa_{\ell}^{S}$
$h$   $\cos\alpha/\sin\beta$   $\cos\alpha/\sin\beta$   $-\sin\alpha/\cos\beta$
$H$   $\sin\alpha/\sin\beta$   $\sin\alpha/\sin\beta$   $\cos\alpha/\cos\beta$
$A$   $\cot\beta$   $-\cot\beta$   $\tan\beta$
Table 1: Yukawa couplings of the fermions $f=u,d$ and $\ell$ to the neutral Higgs bosons $S=h,H$ and $A$ in the 2HDM Type-X.
III Bounds and constraints in the parameter space scans
To determine the regions that meet both theoretical requirements and experimental observations, the subsequent numerical explorations of the potential parameter spaces have been carried out:
$M_{{h}}\in[90,125]~{}(\text{GeV}),\quad M_{{H}}=125~{}(\text{GeV}),\quad M_{{A}}\in[90,125]~{}(\text{GeV})$ (4)
$M_{{H}^{\pm}}\in[90,125]~{}(\text{GeV}),\quad\text{sin}(\beta-\alpha)\in[-0.25,0.05],\quad\text{tan}\beta\in[20,200]$
$m_{12}^{2}=M_{{H}^{\pm}}^{2}\frac{\tan\beta}{1+\tan^{2}\beta}~{}(\text{GeV})^{2}.$
The program 2HDMC-1.8.0 Eriksson:2009ws was employed to calculate the Higgs branching ratios (BRs) and to examine the following theoretical and experimental constraints.
  • Vacuum stability Deshpande:1977rw : the scalar potential should be bounded by the conditions $\lambda_{1}>0$, $\lambda_{2}>0$, $\lambda_{3}>-\sqrt{\lambda_{1}\lambda_{2}}$ and $\lambda_{3}+\lambda_{4}-\lambda_{5}>-\sqrt{\lambda_{1}\lambda_{2}}$.
  • Perturbativity constraints Kanemura:1993hm ; Akeroyd:2000wc , which imply the condition $|\lambda_{i}|\textless 8\pi$, which holds for $i=1$ to 5.
  • Tree-level perturbative unitarity Kanemura:1993hm ; Arhrib:2000is ; Akeroyd:2000wc : requiring all $2\to 2$ scattering processes involving Higgs and gauge bosons at high energy to be unitary.
  • Oblique EW parameters ($S$, $T$ and $U$Peskin:1990zt ; Peskin:1991sw , which govern the mass splitting between the Higgs states, with the following measured values ParticleDataGroup:2022pth :
    $S=-0.02\pm 0.10,\quad T=0.03\pm 0.12.$ (5)
    The correlation factor between $S$ and $T$ is set to $0.92$ for consistency at $95\%$ Confidence Level (CL).
To account for potential additional Higgs bosons, exclusion bounds at $95\%$ CL are enforced using the HiggsBounds-5.9.0 program Bechtle:2020pkv , which systematically checks each parameter point against the $95\%$ CL exclusion limits derived from Higgs boson searches conducted by LEP, Tevatron and LHC experiments.
To ensure agreement with the measurements of the SM-like Higgs state, constraints are enforced using the HiggsSignals-2.6.0 program Bechtle:2020uwn , which incorporates the combined measurements of the SM-like Higgs boson from LHC Run-1 and Run-2 data.
The constraints of flavour physics are incorporated using SuperIso v4.1 Mahmoudi:2008tp using the following observables:
  • BR$(B\to X_{s}\gamma)=(3.32\pm 0.15)\times 10^{-4}$ HFLAV:2022esi ,
  • ${\rm BR}(B^{0}\to\mu^{+}\mu^{-})_{\text{(LHCb)}}$=$\left(1.2^{+0.8}_{-0.7}\right)\times 10^{-10}$ LHCb:2021awg ; LHCb:2021vsc ,
  • ${\rm BR}(B^{0}\to\mu^{+}\mu^{-})_{\text{(CMS)}}$=$\left(0.37^{+0.75}_{-0.67}\right)\times 10^{-10}$ CMS:2022mgd ,
  • ${\rm BR}(B_{s}\to\mu^{+}\mu^{-})_{\text{(LHCb)}}$ = $\left(3.09^{+0.46}_{-0.43}\right)\times 10^{-9}$ LHCb:2021awg ; LHCb:2021vsc ,
  • ${\rm BR}(B_{s}\to\mu^{+}\mu^{-})_{\text{ (CMS)}}$=$\left(3.83^{+0.38}_{-0.36}\right)\times 10^{-9}$ CMS:2022mgd ,
  • BR$(B\to\tau\nu)=(1.09\pm 0.20)\times 10^{-4}$ HFLAV:2022esi .
Figure 1: Scatter plot of ${\rm{BR}}(h\to\tau^{+}\tau^{-})\times{\rm{BR}}(A\to\tau^{+}\tau^{-})$ as a function of $M_{{h}}$ and $M_{{A}}$ (left panel) and $M_{{A}}-M_{{h}}$ as a function of $\sin(\beta-\alpha)$ and $\tan\beta$ (right panel).
[…]
When the constraint on $a_{\mu}$ is considered, there are two types of contributions to $\Delta a_{\mu}$, which are one-loop corrections from $H$, $A$ and $H^{\pm}$ Chun:2016 , as well as two-loop Barr-Zee contributions with heavy fermions Ilisie:2015 . Only in the large $\text{tan}\beta$ regions $\Delta a_{\mu}$ is large enough to explain the newest FNAL measurements. The surviving parameter space is shown in Fig. 2. Compared with Fig. 1 (left), most parameter spaces have been excluded; in fact, only points with $\text{tan}\beta>130$ are available, and the accessible ($M_{h},M_{A}$) regions are rather sparse.
Figure 2: Scatter plot of $\tan\beta$ as a function of $M_{{h}}$ and $M_{{A}}$ when the $a_{\mu}$ effect is considered in additional to the constraints imposed in the previous figure.
To test the allowed parts of the parameter space, nine BPs are given in the Tab. 2, in terms of the 2HDM Type-X parameters entering our production and decay process, including the corresponding $pp\to Z^{*}\to hA\to 4\tau$ cross sections at the LHC with $\sqrt{s}=14$ TeV. Based on this table, we select two scenarios. For BP1-4, where $\tan\beta$ always below 100 and $\sin(\beta-\alpha)$ is close to but not exactly zero. Here, $a_{\mu}$ remains a concern (till new measurements are performed) For BP5-9, the model is in the alignment-limit scenario, where $\sin(\beta-\alpha)=0$ and $\tan\beta$ is very large. In this scenario, $a_{\mu}$ can be explained well (with the present measurements).
$M_{h}$ $M_{A}$ $M_{H}^{\pm}$ $\tan\beta$ $\sin(\beta-\alpha)$ $\sigma_{4\tau,~{}14~{}{\rm TeV}}$ (fb)
BP1 95.10 121.70 95.02 24.19 $-0.076$ 38.76
BP2 120.10 118.50 119.80 41.25 $-0.003$ 27.89
BP3 99.52 102.30 99.53 60.71 $-0.034$ 52.01
BP4 109.10 89.07 109.00 84.43 $-0.025$ 55.23
BP5 105.70 117.40 105.60 105.30 $-0.001$ 35.62
BP6 113.00 103.20 113.00 137.80 0.0 39.79
BP7 109.70 115.30 109.70 143.00 0.0 34.24
BP8 107.00 85.31 107.00 166.10 0.0 60.71
BP9 116.20 111.20 116.20 196.70 0.0 31.79
Table 2: 2HDM Type-X input parameters and Leading Order (LO) cross sections (at the parton level with $\sqrt{s}=14$ TeV) are presented for each BP. The unit of all masses is GeV and we fix the SM-like Higgs boson mass as $M_{H}$ = 125 GeV. (Also recall that $m_{12}$ has been taken as a derived quantity, function of $M_{H^{\pm}}$ and $\tan\beta$.)
IV Collider phenomenology
In this section, we present a detailed MC analysis for both signal and background events at the detector level. In the $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$ process, i.e., the light Higgs (pseudo)scalar decays into two $\tau$ leptons and one of the $\tau$’s decays into a charged lepton and neutrino, while the other $\tau$ decays hadronically (thus labeled as $\tau_{h}$).
The main SM backgrounds are: $t\overline{t}\to l\nu_{l}l\nu_{l}b\overline{b}$, $W^{\pm}tb\to l\nu_{l}l\nu_{l}b\overline{b}$ (where the $W^{\pm}$ boson and the $b$ jet are not from a top quark resonance), $WWjj\to l\nu_{l}l\nu_{l}jj$, $Zjj\to l\nu_{l}l\nu_{l}jj$, $ZZ\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$, $t\overline{t}Z\to l\nu_{l}l\nu_{l}b\overline{b}\tau\tau$, $t\overline{t}ZZ\to l\nu_{l}l\nu_{l}l\nu_{l}b\overline{b}\tau\tau$ and $t\overline{t}WW\to l\nu_{l}l\nu_{l}l\nu_{l}jj$. To suppress the huge SM background process $Zjj$, we deliberately choose two SS leptons plus two hadronic $\tau$’s (for signal events, this is about $10\%$ of the total number of $4\tau$ ones). As a result, the final state consists of two SS leptons and two (light) jets.
IV.1 Event generation and selection
We use MadGraph5-aMC@NLO v3.4.0 Alwall:2014hca to calculate cross sections and generate both signal and background events at the parton level. The following acceptance cuts are adopted for the signal and background generations:
$|\eta(l,j)|<2.5,~{}p_{\text{T}}(l,j)>10~{}\text{GeV},~{}E_{T}^{\text{miss}}>10~{}\text{GeV},~{}\Delta R(ll,lj,jj)>0.4.$ (6)
The parton-level events are then passed to Pythia Sjostrand:2006za ; Sjostrand:2014zea to simulate initial- and final-state radiation (i.e. the QED and QCD emissions), parton shower, hadronisation and heavy flavour decays. We further use Delphes v3.4.2 deFavereau:2013fsa to simulate the detector effects. For each event, the anti-$k_{t}$ jet algorithm Cacciari:2008gp is used to cluster the jets with the jet parameter $\Delta R=0.4$ in the FastJet package Cacciari:2011ma . The following kinematic cuts are further adopted in order to emulate the detector acceptance:
$|\eta(l,j)|<2.5,~{}p_{T}(l,j)>10~{}\text{GeV}.$ (7)
The leptons are used for triggering purposes, so their kinematics must satisfy the trigger requirements [^1] The CMS collaboration has developed a double-muon trigger to search for low $P_{T}$ muons from $B$ meson decays that are close together. A similar approach has been adopted recently to search for electron pairs from $B$ meson decays, by tightening the topological selection criteria on the two Level-1 electron objects. In our recent paper Arhrib:2023apw , we explored the possibility of also developing an electron-muon trigger to search for low $P_{T}$ leptons ($P_{T}\sim 10$ GeV)..
Then, we select a final state consisting of two leptons and two jets from detector-simulated events. Tau tagging is not applied to improve the signal event rate. The signal results are presented in Tab. 3 and the relevant background processes are summarised in Tab. 4. The number of background events appears to be approximately $10^{5}$ larger than the number of signal events, with the dominant background processes at this stage $pp\to t{\bar{t}}$ and $pp\to Zjj$.
To further suppress the huge SM backgrounds, we impose the selection condition requiring SS leptons. After applying this criterion, the number of signal and background events becomes comparable, with the dominant background contributions arising from the $pp\to t{\bar{t}}$ and $pp\to ZZ$ processes.
The cross sections of the signal and background processes after the detector acceptance cuts, the selection condition assuming a $2l2j$ final state, and plus the two SS leptons requirement are shown in Tabs. 3 and  4, respectively.
$\sigma$ (fb) BP1 BP2 BP3 BP4 BP5 BP6 BP7 BP8 BP9
Parton level 10.96 8.75 13.66 14.16 10.43 11.24 10.09 14.79 9.49
Selection of $2l2j$ 3.02 2.59 3.66 3.68 2.87 3.20 2.95 3.83 2.67
Selection of SS $2l$ 1.196 1.494 1.522 1.498 1.007 1.239 1.141 1.301 1.295
Table 3: Cross sections for all signal BPs at the parton level as well as the detector level after selection cuts at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$.
$\sigma$ (fb) $t\bar{t}$ $W^{\pm}tb$ $W^{+}W^{-}jj$ $Zjj$ $ZZ$ $t\bar{t}Z$ $t\bar{t}ZZ$ $t\bar{t}W^{+}W^{-}$
Parton level 16060 518.3 1053 317600 18.89 0.49 $1.14\times 10^{-4}$ 0.02
Selection of $2l2j$ 8787.7 289.9 530.1 151086 10.0 0.33 $1.1\times 10^{-4}$ 0.018
Selection of SS $2l$ 19.43 0.62 1.99 0 2.51 0.079 $3.3\times 10^{-5}$ $7.6\times 10^{-3}$
Table 4: Cross sections for all backgrounds at the parton level as well as the detector level after selection cuts at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$.
In Fig. 3, the transverse momentum distributions of the (second-)leading jets (frame (a)) and leptons (frame (b)) are presented for the signal (e.g. BP4) and all background events. It is clear that the transverse momenta of the signal final state are always much softer than those for background events: This is because they emerge from rather light Higgs bosons, whose energies are capped at masses of the order of 100 GeV.
Figure 3: Normalised distributions in transverse momentum of the first- and second-leading jets (a) and leptons (b) for BP4 and backgrounds at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$.
IV.2 Event reconstruction
To further improve the signal-to-background discrimination, we reconstruct some kinematic features of signal events. However, it is extremely difficult to completely reconstruct the momentum of each light Higgs state. This is because each Higgs boson decays into two $\tau$, with one of the $\tau$’s decaying to a lepton and two neutrinos plus two $\tau$ neutrinos emerging from hadronic $\tau$ decays, leading to the presence of six neutrinos in a signal event.
Nevertheless, since the lepton momenta and the missing $E_{T}$ (MET), where $E_{T}$ refers to transverse energy, are all very small, we can approximately reconstruct the light Higgs bosons with only one lepton and one jet. Thus, there are two possible combinations with two leptons and two $\tau$ jets in each event. By defining a suitable $\chi^{2}$,
$\chi^{2}={(M_{{lj}}^{1}-M_{{h}})}^{2}+{(M_{lj}^{2}-M_{A})}^{2},$ (8)
where $M_{h(A)}$ is the mass of the $h$($A$) (pseudo)scalar. We pair the two leptons and two jets to find a combination that minimises it and then assign these to $M_{lj}^{1}$ and $M_{lj}^{2}$, respectively. Since we did not consider neutrinos during such a mass reconstruction, the final $M_{lj}^{1,2}$ values should be smaller than the real $M_{h/A}$ ones; see Figs. 4 frames (a) and (b). Taking again BP4 as an example, the reconstructed light Higgs mass distributions are shown in Fig. 4. A weak positive correlation was observed between these two variables (with $r$ value of 0.27), as exemplified in Fig. 4 frame (c).
Figure 4: Distributions in invariant mass of a lepton and a jet for BP4 and backgrounds at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$. The reconstructed values, corresponding to the best $\chi^{2}$ fit, are labeled as $M_{lj}^{1}$ and $M_{lj}^{2}$, which are normalised in (a) and (b), respectively, and unnormalised (for $L~{}=~{}\text{300}~{}\text{fb}^{-1}$) when correlated in (c). In the bottom frame the signal (background) is in blue(red).
In order to further separate the signal from background events, we can also combine the two leptons and two jets, denoting their invariant masses as $M_{ll}$ and $M_{jj}$, respectively. For some SM background processes, these lepton and/or jet pairs always come from some heavy resonance, i.e.. a top (anti)quark and a $W^{\pm}$ or $Z$ boson, the background spectra should have peaks in special mass regions. However, such a feature is largely removed when selecting SS leptons, as shown in Fig. 5, so it cannot really be exploited here.
Figure 5: Normalised distributions in invariant mass of lepton (a) and jet (b) pairs for BP4 and backgrounds at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$.
Figure 6: Normalised distributions in visible invariant mass (a) and $H_{T}$ (b) for BP4 and backgrounds at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$.
Figure 7: Normalised distributions in missing transverse energy (a) and $M_{\text{T2}}$ (b) for BP4 and backgrounds at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$. In the right frame we show the three classes of background identified in the text separately.
We can further reconstruct the visible invariant mass, i.e. the one made up of the two leptons and two jets, $M_{\rm vis}$, but without any MET. In the signal, these particles come from a virtual resonance $Z^{*}$, hence $M_{\rm vis}$ should be smaller than in the backgrounds, where this is not normally the case, the latter thus yelding a larger value; see Fig. 6 frame (a). Another useful kinematic variable is $H_{T}$, which is the scalar sum of the transverse momenta of the final-state objects, shown in Fig. 6 frame (b). These two observables are generated in a very similar way between signal and background, which is evident by comparing the two frames in this figure.
In our process, two neutrinos come from leptonic $\tau$ decays, hence very soft because the $\tau$ leptons are from cascade decays. There are also neutrinos in the backgrounds from $W^{\pm}$, $Z$ boson or top (anti)quark (semi)leptonic decays, which have momenta higher than those of the neutrinos in signal events. The MET distribution is shown in Fig. 7 (a). Furthermore, the $M_{\text{T2}}$ variable is often used to track the masses of unseen pairs of particles Lester:1999ma ; Barr:2003ma , which is defined as:
[…]
We thus reconstruct the $M_{\text{T2}}$ variable for both signal and background, while splitting the latter into three classes: top (including $t\bar{t}$, $t\bar{t}W^{+}W^{-}$, $t\bar{t}Z$, $t\bar{t}ZZ$ events), $W^{\pm}$ boson (including $W^{+}W^{-}jj$, $W^{\pm}tb$ events) and $Z$ boson (including $ZZ$ events). The $M_{\text{T2}}$ distributions are shown in Fig. 7 (b). It can be seen that those of the signal and $Z$ boson background are below 100 GeV or so while they are close to 150 GeV or so for the top quark and $W^{\pm}$ boson processes. Thus, the $M_{\text{T2}}$ variable provides an efficient discrimination, at least, against the top quark and $W^{\pm}$ boson noises.
IV.3 Multivariate analysis
To optimise the signal and background distinction, we exploit a ML analysis, that is, a Gradient-Boosted Decision Tree (GBDT) approach is further applied, which is implemented in the Toolkit for Multivariate Date Analysis (TMVA) with ROOT Therhaag:2010zz .
A total of ten input variables are used in such a GBDT/TMVA analysis, as listed in Tab. 5. In addition to the invariant masses mentioned above, $M_{\text{T2}}$, and $H_{T}$, we also compute five angular variables between pairs of final-state objects: $\cos(\theta_{l_{1}j_{1}})$, $\cos(\theta_{l_{1}j_{2}})$, $\cos(\theta_{l_{2}j_{1}})$, $\cos(\theta_{l_{2}j_{2}})$. These angles are defined as the cosine of the angle between the two final states leptons $l$ and jets $j$, ordered in $E_{T}$, and $\cos(\theta_{lj-lj})$ is the angle between two reconstructed bound states $(lj)^{1}$ and $(lj)^{2}$. The distributions are shown in Fig. 8. As seen here, these variables are expected to be particularly useful since, in the signal process, they tend to be collinear due to the underlying boost structure, unlike in the background case.
Energy variables $M_{lj}^{1}$ $M_{lj}^{2}$ $M_{ll}$ $M_{jj}$ $H_{T}$ $M_{\text{T2}}$
Angular variables $\cos(\theta_{lj-lj})$ $\cos(\theta_{l_{1}j_{1}})$ $\cos(\theta_{l_{1}j_{2}})$ $\cos(\theta_{l_{2}j_{1}})$ $\cos(\theta_{l_{2}j_{2}})$
Table 5: The input observables used in the GBDT analysis.
Figure 8: Normalised distributions in the (cosine of the) angle between pairs of jets/leptons (a)-(d), and two reconstructed bound state $(lj)^{1}$ and $(lj)^{2}$ (e) for BP4 and backgrounds at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$.
In the training stage, we take the signal and $ZZ$ background events as input. Because their physical process are very similar and difficult to distinguish apart in the selection with kinematic cuts. The $M_{\text{T2}}$ and the invariant mass $M_{lj}^{1,2}$ are the most important variables while training. But the angular variables $\cos(\theta_{lj-lj})$ and $\cos(\theta_{l_{i}j_{k}}),~{}i\neq k$ are also useful. Ultimately, for the GBDT output, a very good separation between the signal and background can be achieved, which is shown in Fig. 9.
Figure 9: GBDT output for BP4 and backgrounds.
IV.4 Significances
The process of searching for signals at high-energy colliders using ML techniques is ultimately a statistical process. Therefore, to look for a signal, a large number of events need to be collected, and the kinematic variables of the final-state particles need to be analysed to find differences from the noise, but then a sizeable statistical sample needs to be collected, so as to have enough significance, in order to claim sensitivity to the signal which is being searched for.
The latter is defined here as $\Sigma=\frac{N_{S}}{\sqrt{(N_{S}+N_{B})}}$, where $N_{S}$ and $N_{B}$ represent the number of events for the signal ($B$) and (total) background ($B$), respectively, at a given luminosity $L$.
To distinguish the signal from the backgrounds and to compute the ensuing significance, we apply six kinematic cuts on reconstructed observables and one cut on the GBDT output, the details of which are listed in Tab. 6 (for BP4 in the signal case) and where $L~{}=~{}\text{300}~{}\text{fb}^{-1}$ (with a centre-of-mass energy of 14 TeV, as usual). Following the acceptance cuts and upon applying the SS lepton criterion (the combination of which we refer to as preselection, hereafter), we see a strong decrease of the initially huge SM backgrounds to the same order as the signal. For example, the troublesome $Zjj$ is almost eliminated after this cut (thus, it is no longer included in this table). As Figs. 46 show, the kinematic selection is already quite efficient in separating the signal from the background, especially for the variable $M_{\text{T2}}$. Finally, we also applied a GBDT discrimination before computing the final significances.
$L~{}=~{}\text{300}~{}\text{fb}^{-1}$ $S$ Top quark $W^{\pm}$ boson $Z$ boson Total background $S/B$ $\Sigma$
Acceptance 938.2 $1.10\times 10^{6}$ 60544.8 2584.9 $1.16\times 10^{6}$ 0.008 0.87
SS Leptons 514.8 2970.5 301.4 698.6 3970.4 0.130 7.69
$HT\in[60,400]~{}\text{GeV}$ 492.8 2743.7 260.6 682.3 3686.5 0.134 7.62
$M_{ll}\in[10,300]~{}\text{GeV}$ 490.7 2548.3 235 675.7 3459 0.142 7.81
$M_{lj}^{1}\in[0,120]~{}\text{GeV}$ 434.8 1656.5 139.9 486.5 2282.9 0.190 8.34
$M_{lj}^{2}\in[0,150]~{}\text{GeV}$ 415.6 1490.1 123.1 457.6 2070.8 0.20 8.34
$M_{vis}\in[80,500]~{}\text{GeV}$ 364.7 950.5 88.7 393.1 1432.3 0.255 8.6
$MT^{2}\in[10,110]~{}\text{GeV}$ 293.8 0 0 149.5 149.5 1.965 13.96
${\rm GBDT}\in[-0.35,1]$ 247.8 0 0 39.6 56.8 4.363 14.2
Table 6: Response to our selection cuts for the signal (BP4) and background (separately and total) rates computed at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$.
The results on $\Sigma$ for all 9 BPs are shown in Tab. 7. With only the acceptance and SS lepton criteria, the significances can reach 3. After applying the kinematic cuts, including exploiting the $M_{\text{T2}}$ observable, the significances can be even larger than 10. Upon finally applying the GBDT selection, $\Sigma$ values can increase even further (albeit only slightly). In short, it is clear that a judicious combination of preselection and kinematic cuts combined with a GBDT analysis can result in large sensitivity to a significant expanse of 2HDM Type-X parameter space through our chosen signal at the current LHC stage, already.
$\Sigma$ BP1 BP2 BP3 BP4 BP5 BP6 BP7 BP8 BP9
After selecting SS leptons 6.22 5.11 7.71 7.69 5.94 6.42 5.88 8.08 5.4
After kinematic cuts, w/o $M_{\text{T2}}$ GBDT 8.31 7.46 10.71 10.14 8.15 7.91 7.93 10.87 7.78
With $M_{\text{T2}}$, w/o GBDT 11.76 10.81 14.38 13.96 11.54 11.33 11.37 14.65 11.24
With $M_{\text{T2}}$ GBDT 11.91 10.85 14.23 14.20 11.94 11.57 11.52 14.31 11.12
Table 7: Significances following our different selections for all signals (BP1-9) at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$.
V Conclusions
In this paper, we have investigated the feasibility of hadro-producing the light Higgs scalar $h$ and pseudoscalar $A$ through $Z^{*}$ exchange (initiated by $q\bar{q}$ pairs) at the LHC in the 2HDM Type-X (or lepton specific). We have focused on the $4\tau$ final state, where two $\tau$’s decay leptonically to form a pair of SS leptons (electrons and/or muons) and two $\tau$’s decay hadronically (thus producing two jets). We have proven that this kind of signal could be found (or excluded) at the end of Run 3 of the LHC, following a dedicated selection based on both a kinematic and GBDT/TMVA analysis. In order to motivate future studies of this signature, we have also proposed 9 BPs and processed them through a realistic MC analysis for a centre-of-mass energy of $14$ TeV and (integrated) luminosity $L~{}=~{}\text{300}~{}\text{fb}^{-1}$, thus emulating actual LHC conditions.
Acknowledgments
SM is supported in part through the NExT Institute and the STFC Consolidated Grant ST/X000583/1. SS is supported in full through the NExT Institute and acknowledges the use of the IRIDIS High Performance Computing Facility, and associated support services, at the University of Southampton, in the completion of this work. YW’s work is supported by the Natural Science Foundation of China Grant No. 12275143, and the Fundamental Research Funds for the Inner Mongolia Normal University Grant No. 2022JBQN080. QSY is supported by the Natural Science Foundation of China under the Grants No. 11875260 and No. 12275143.